OpenStudy (anonymous):

A grinding wheel (which is cylindrical) has radius 4.0 cm and a mass of 800 g. a. Find the wheel’s moment of inertia. b. What net torque must be applied to achieve an angular acceleration of 150 rad/s2? c. With this acceleration, how many revolutions would the wheel make in getting up to a speed of 600 revolutions per minute?

OpenStudy (experimentx):

One question at time: a) http://upload.wikimedia.org/wikipedia/en/math/2/9/b/29bac02573ea5d0bbf08a7506e4e9b37.png Use Iz, that is moment of inertia of circular disc

OpenStudy (experimentx):

b) Torque = Inertia x angular acceleration

OpenStudy (anonymous):

so a) = 0.000512 kgm^2

OpenStudy (experimentx):

haven't calculated yet ... possibly. seems that you changed in SI

OpenStudy (anonymous):

yes they like our answers in kg and m

OpenStudy (experimentx):

for c) \( (\omega_2)^2 = (\omega_1)^2 + 2 \alpha\theta\) where omega2 is 600rev per min, omega2 = 0, alpha is angular accn. find theta, and divide it by 2pi // that will be your answer.

OpenStudy (experimentx):

I like in kg and m too.

OpenStudy (anonymous):

b is .0768 Nm?

OpenStudy (anonymous):

for c i got 190.985 revolutions

OpenStudy (experimentx):

not really sure ... if your a) is correct then b) must be 150/pi * 0.000512

OpenStudy (experimentx):

b) must be 150/2pi * 0.000512

OpenStudy (experimentx):

c) must be (600 *2pi/60)^2 = 2 * (150/2pi) theta divide theta by 2pi, you will get rev.

OpenStudy (anonymous):

b is .012223 then ?

OpenStudy (experimentx):

I haven't calculated .. I am not quite sure about changing radian/sec like that in angular acceleration.

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Radian_per_second

OpenStudy (anonymous):

im getting a little over 13 revolutions. im not sure if that seems like a legitimate answer or not

OpenStudy (experimentx):

you don't have answer at the back of your text??

OpenStudy (anonymous):

no its questions that were written for us. Not in a text book

OpenStudy (experimentx):

then best would be to search similar questions that have answers /// usually my answers are incorrect.

OpenStudy (experimentx):

***numerical answers.